OpenStudy (anonymous):

sold catch of 184 lemon and pepper fish for 1506.20. Lemonfish sold for 13.95 per pound and pepper fish for 6.25 per pound. How much of each were sold

OpenStudy (amistre64):

L + P = 184 L = 184-P L(13.95) + P(6.25) = 1506.20 got it so far?

OpenStudy (amistre64):

(184 -P)(13.95) + P(6.25) = 1506.20 2566.8 - 13.95P + 6.25P = 1506.20 you still with me?

OpenStudy (amistre64):

2566.80 -1506.20 = 13.95P - 6.25P 1060.60 = P(13.95 - 6.25) you keeping up with me :)

OpenStudy (amistre64):

1060.60 = P(7.70) 1060.60/(7.70) = P P = 137.74

OpenStudy (amistre64):

L = 184 - P L = 184 - 137.74 = 46.26

OpenStudy (anonymous):

Hey amistre can you help me on my problem please?

OpenStudy (amistre64):

sure, I think I handled this one good :)

OpenStudy (anonymous):

its at the very top

OpenStudy (amistre64):

rancher one?

OpenStudy (anonymous):

yea